The perimeter of a right angled triangle having 24 cm base is 60 cm.find out the remaining two sides of the triangle.

Answer by richwmiller(17219)

60-24=36
either
a+b=36,
a^2+b^2=24^2
24 is not the hypotenuse.
So
a+b=36,
a^2+24^2=b^2
a = 10, b = 26